A car traveling at 33 m/s slows to 7.5 m/s while driving 410 m. Calculate the car's acceleration.


1
Expert's answer
2021-09-22T07:05:56-0400
"v_f^2-v_i^2=2as\\7.5^2-33^2=2(410)a\\a=-1.3\frac{m}{s}"